After spending the day at Great America Mike heads north traveling 55 mph and John heads south traveling 60 mph. How long will it take for Mike and John to be 17.25 miles apart? ( state the time in minutes)
Answer by Alan3354(69443)

After spending the day at Great America Mike heads north traveling 55 mph and John heads south traveling 60 mph. How long will it take for Mike and John to be 17.25 miles apart? ( state the time in minutes)
-----------
Going in opposite means their speeds are added.
55 + 60 = 115 mph
17.25/115 = 69/460 = 3/20 hours
= 9 minutes